Miller Company's total sales are $153,000. The company's direct labor cost is $18,360, which represents 30% of its total conversion cost and 40% of its total prime cost. Its total selling and administrative expense is $22,950 and its only variable selling and administrative expense is a sales commission of 5% of sales. The company maintains no beginning or ending inventories and its manufacturing overhead costs are entirely fixed costs.


 Required:

 1. What is the total manufacturing overhead cost?

 2. What is the total direct materials cost?

 3. What is the total manufacturing cost?

 4. What is the total variable selling and administrative cost?

 5. What is the total variable cost?

 6. What is the total fixed cost?

 7. What is the total contribution margin?
Answer #1

1. Conversion cost = direct labor+manufacturing overhead

Now 30% of conversion cost = direct labor cost. Thus 30% of conversion cost = 18360

Thus conversion cost = 18360/30% = $61200

Now, Conversion cost = direct labor+manufacturing overhead. Thus 61200 = 18360+manufacturing overhead

or manufacturing overhead = 61200-18360 = $42840

2. Total prime cost = direct labor cost/40% = 18360/40% = 45,900

Total manufacturing cost = prime cost+manufacturing overhead cost. Total manufacturing cost = direct materials+direct labor+manufacturing overhead.

Thus direct materials+direct labor+manufacturing overhead = prime cost+manufacturing overhead

or direct materials+18360+42840= 45,900+42840

or direct materials cost = 45900-18360 = $27540

3. Total manufacturing cost = direct material+direct labor+manufacturing overhead = 27540+18360+42840= $88740

4. Variable selling and administrative cost = 5% of sales = 5% of 153,000 = $7,650

5. Total variable cost = $7650+direct materials+direct labor = 7650+27540+18360 = $53,550

6. Total costs = total manufacturing cost+total selling and administrative expense = 88740+22950 = $111,690

Total fixed cost = total cost-total variable cost = 111690-53550 = $58140

7. Contribution margin = sales - variable costs = 153,000-53550

= $99,450